# CRASAR-U-DROIDs: A Large Scale Benchmark Dataset for Building Alignment and Damage Assessment in Georectified sUAS Imagery

### v1

@@ II Prior Work
A review of literature identifies five sUAS datasets specifically from disasters, but have gaps which are addressed by CRASAR-U-DROIDs: they are not georectified; represent a much smaller scale of spatial area (67.98km2 vs approximately 4.17km2), pixels (67.035 gigapixels vs 53.99 gigapixels), represent a smaller number of disasters (10 vs 6), and fewer types of disasters (5 vs 2); have less utility for emergency management because the labeling schemas were ad hoc or a misapplication of an existing standard; and do not consider spatial alignment. These five datasets are a subset of a total of eleven existing labeled disaster imagery datasets from any aerial asset (4 unmanned, 2 manned, 4 satellite, and 1 both manned and unmanned) identified via a systematic search within the IEEE Xplore, ACM Digital Library, Science Direct scientific databases, and the natural hazards DesignSafe Data Depot [1]. The criteria for choosing the 11 datasets was to include only those with i) aerial imagery from disasters with significant economic consequences which overwhelmed local response capabilities, consistent with the definition of a disaster, ii) labeled damage to populated areas and structures rather than attributes such as boundaries of affected areas, in keeping with the CRASAR-U-DROIDs focus on building damage assessment iii) the dataset was at least partially available for download. Seventeen datasets failed to meet the criteria for inclusion: eight fires [56, 57, 52, 61, 12, 59, 19, 16], four floods [41, 10, 55, 44], one landslide [62], one earthquake [13], and three which included multiple types of events [30, 60, 34]. Six datasets of labeled georectified data collected from sUAS that are not explicitly from disasters [50, 62, 14, 6, 63, 43] are included in this review as a separate discussion on characterizing work on labeled georectified data, as such data is a key element of CRASAR-U-DROIDs.
@@ II-A Aerial Imagery Disaster Response Datasets
The 11 aerial imagery datasets for disaster response and assessment are divided into three categories based on the sensing platform that collected the imagery and detailed below: sUAS (Sec. II-A1), manned aircraft (Sec. II-A2), and satellite imagery (Sec. II-A3). Each dataset is described in terms of the source event, the annotation schema, and the key attributes (area, tasks, etc.), which are summarized in Table II, with Tables III and IV providing additional details on the five sUAS disaster datasets. As Volan v.2018 [48] contains imagery from multiple types of sensing platforms, it is discussed in multiple subsections.
Table II is the overarching comparison of all 11 datasets grouped by sensing platform and supports the claims that CRASAR-U-DROIDs is the only georectified sUAS disaster dataset and the largest in terms of spatial area and number of pixels. In order to facilitate the comparison, the key attributes (columns 2-8) merit further explanation:
@@ II-A2 Manned Aircraft Disaster Imagery Datasets
The (Chen et. al. 2018) dataset [15, 18] contains bounding box and building polygon annotations of building damage for areas impacted by Hurricane Harvey. The publication claims their annotations represent a combination of annotations from crowd workers and from FEMA flood damage model estimates as well as both satellite and manned aerial imagery[15]. However, only annotations for building damage from overflights of NOAA Manned Aircraft are available and referenced online [18]. Further, while annotations could be retrieved for 566,659 buildings, they do not appear to contain the crowd-sourced annotations that are described in the publication. Instead, only labels that correspond to the FEMA flood damage model estimates. All available bounding boxes are annotated as either “none”, “AFF,” “MIN,” “MAJ,” or “DES”. As a result, it is difficult to accurately assess the size and relevance of this dataset as it is not fully available. Due to this lack of certainty, this dataset has been omitted from Table II.

### v2

@@ II Prior Work
A review of literature identifies five sUAS datasets specifically from disasters, but have gaps which are addressed by CRASAR-U-DROIDs: they are not georectified; represent a much smaller scale of spatial area (67.98km2 vs approximately 4.17km2), pixels (67.13 gigapixels vs 53.99 gigapixels), represent a smaller number of disasters (10 vs 6), and fewer types of disasters (5 vs 2); have less utility for emergency management because the labeling schemas were ad hoc or a misapplication of an existing standard; and do not consider spatial alignment. These five datasets are a subset of a total of twelve existing labeled disaster imagery datasets from any aerial asset (4 unmanned, 3 manned, 4 satellite, and 1 both manned and unmanned) identified via a systematic search within the IEEE Xplore, ACM Digital Library, Science Direct scientific databases, and the natural hazards DesignSafe Data Depot [1]. The criteria for choosing the 12 datasets was to include only those with i) aerial imagery from disasters with significant economic consequences which overwhelmed local response capabilities, consistent with the definition of a disaster, ii) labeled damage to populated areas and structures rather than attributes such as boundaries of affected areas, in keeping with the CRASAR-U-DROIDs focus on building damage assessment iii) the dataset was at least partially available for download. Seventeen datasets failed to meet the criteria for inclusion: eight fires [56, 57, 52, 61, 12, 59, 19, 16], four floods [41, 10, 55, 44], one landslide [62], one earthquake [13], and three which included multiple types of events [30, 60, 34]. Six datasets of labeled georectified data collected from sUAS that are not explicitly from disasters [50, 62, 14, 6, 63, 43] are included in this review as a separate discussion on characterizing work on labeled georectified data, as such data is a key element of CRASAR-U-DROIDs.
@@ II-A Aerial Imagery Disaster Response Datasets
The 12 aerial imagery datasets for disaster response and assessment are divided into three categories based on the sensing platform that collected the imagery and detailed below: sUAS (Sec. II-A1), manned aircraft (Sec. II-A2), and satellite imagery (Sec. II-A3). Each dataset is described in terms of the source event, the annotation schema, and the key attributes (area, tasks, etc.), which are summarized in Table II, with Tables III and IV providing additional details on the five sUAS disaster datasets. As Volan v.2018 [48] contains imagery from multiple types of sensing platforms, it is discussed in multiple subsections.
Table II is the overarching comparison of 11 of the 12 datasets grouped by sensing platform and supports the claims that CRASAR-U-DROIDs is the only georectified sUAS disaster dataset and the largest in terms of spatial area and number of pixels. In order to facilitate the comparison, the key attributes (columns 2-8) merit further explanation:
@@ II-A2 Manned Aircraft Disaster Imagery Datasets
The (Chen et. al. 2018) dataset [15, 18] contains bounding box and building polygon annotations of building damage for areas impacted by Hurricane Harvey. The publication claims their annotations represent a combination of annotations from crowd workers and from FEMA flood damage model estimates as well as both satellite and manned aerial imagery[15]. However, only annotations for building damage from overflights of NOAA Manned Aircraft are available and referenced online [18]. Further, while annotations could be retrieved for 566,659 buildings, they do not appear to contain the crowd-sourced annotations that are described in the publication. Instead, only labels that correspond to the FEMA flood damage model estimates are available. All bounding boxes are annotated as either “none”, “AFF,” “MIN,” “MAJ,” or “DES”. As a result, it is difficult to accurately assess the size and relevance of this dataset as it is not fully available. Due to this lack of certainty, this dataset has been omitted from Table II.